OpenStudy (anonymous):

oh my god I hate math! Help please!!! ABCD is a rectangle. find the length of each diagonal. AC=2(5a+1) BD=2(a+1)

OpenStudy (anonymous):

If ABCD is a rectangle, then AC=BD 2(5a+1)=2(a+1) 10a+2=2a+2 8a=0 a=0

OpenStudy (anonymous):

10*0+2 = 2 2*0+2 = 2 Conclusion: the length is 2 units.